y^2+6y+34=0 solve by "completing the square"
Answer by stanbon(75887)

y^2+6y+34=0 solve by "completing the square"
---
y^2+6y + ? = -34 + ?
---
y^2+6y+9 = -34+9
----
(y+3)^2 = -25
----
y + 3 = 5i or y+3 = -5i
-----
y = -3+5i or y = -3-5i
===========================
cheers,
stan H.